OpenStudy (muzzack):

Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 3x - 2y = 15 6x + 5y = 11 4x - 3y = 9 -8x + 6y = -18 5x - 3y = 16 -2x + 3y = -7

OpenStudy (muzzack):

@austinL @annas @AnImEfReaK @aprilann143 @AuroraB @asatt32 @shanerocks @Secret-Ninja @sheelo_mughal @SarahEZZMcK @sparki1 @FMA @FrankiesaysRelax @FatboyMatrix @furrywalls @fire180 @Ria23 @revvy @RosieF @RosieF

OpenStudy (anonymous):

OK IF I DONT KNOW U DONT TAG ME PLZZZ

OpenStudy (anonymous):

what up?

OpenStudy (anonymous):

An infinite number of solutions for a pair of equations would mean both lines run along each other- like the equations are the same or have the same ratios.

OpenStudy (anonymous):

The hint is in the last part.

OpenStudy (muzzack):

thx @AnImEfReaK

OpenStudy (anonymous):

Do you know what the answer might be?

OpenStudy (anonymous):

hold on restate it and explain

OpenStudy (anonymous):

?

OpenStudy (anonymous):

4x -3y = 9 -8 + 6y = -18 has infinitely many solutions

OpenStudy (muzzack):

well the first one is not the answer because it gives 24 = 42

OpenStudy (anonymous):

infinite*

OpenStudy (muzzack):

oh thx :) @FMA

OpenStudy (muzzack):

BECAUSE THE COMMON FACTOR IS 12 RIGHT AND YOU REPLACE IT IN THE OTHER QUESTIONS AND IT GIVE THE SAME NUMBER RIGHT?

OpenStudy (anonymous):

How is the common factor 12?

OpenStudy (muzzack):

3 x 4 = 12 and 6 x 2 = 12 12 is the common factor @AnImEfReaK

OpenStudy (anonymous):

Yes, but where did you get the 6 and 2?